Janet invested $26,000 part at 6% and part at 3%, if the total interest at the end of the year is $1,080, how much did she invest at 6%?
Answer by mananth(16949)

$26,000 total investment
6% ------------$x
3%, ------------26000-x
interest =$1,080,
...
0.06x+0.03(26000-x)=1080
multiply by 100
6x +3(26000-x)=108000
6x+78000-3x=108000
3x=108000-78000
3x=30000
x= $10000 @ 6%
$16000 @ 3%